OpenStudy (anonymous):

Using prime factorization, find the value of each of the following numbers 36 squared

OpenStudy (anonymous):

\[ 36 = 2^2 3^2 \]

OpenStudy (anonymous):

\[ 36^2 = 2^4 3^4=16 (81)=1296 \]